OpenStudy (anonymous):

What is the area of a square with a diagonal of 8 cm

OpenStudy (anonymous):



OpenStudy (anonymous):

use trig to solve for the top and bottom then use the b*h = Area to solve

OpenStudy (anonymous):

Knowing that a square all sides are the same length thus once you find the length of one side all other will be equal, this would also be a right angle triangle

OpenStudy (sandra):

Yes this can be solved with geometry/algebra. What do you know about diagonals of triangles with two equal sides? [use pythagorean theorem]

OpenStudy (anonymous):

Let side of square be = x cm Using Pythagoras Theorem, Diagonal^2 = side^2 + side^2 means 8 = x^2 + x^2 8 = 2x^2 4 = x^2 2 = x so side of square is 2 cm so its area = side*side = 2 cm * 2 cm = 4 cm^2